OpenStudy (anonymous):

how the distributive property relates addition and multiplication. give numerical examples

OpenStudy (anonymous):

They just want you to give an example of the distributive property in action (i.e. using numbers). Thus, as an example, 8( 3 + 7) = 8*3 + 8*7 = 24 + 56 = 80 And that should make sense because you could also simplify it as: 8(3+7) = 8* 10 = 80 Is this helping??

OpenStudy (anonymous):

yes